momentum

a 5 kg fish swimming 1 meter per second swallows an absentminded 1 kg fish swimming toward it at a velocity that brings both to a halt immediately after lunch. what isthe velocity of the smaller fish before lunch?
Answer #1
m1*v1 =m2*v2Therefore, 5kg*1m/s = 1kg * v2The absent minded fish must be swimming at v2 = 5m/s
answered by: BENNETTA
Answer #2
momentum is mass times velocity so in order to have two objectswith the same momentum


5kg * 1m/s=1kg*xm/s

it seems that the answer is 5m/s

answered by: krisiti